I have problems solving this problem could you help me please I am not sure How to do it
Identify the rate, base, and amount in the following applications. Do not solve the applications at this point
1.Paul invested $5000 in a time deposit. What interest will he earn for 1 year if the interest rate is 6.5%?
2.A chemist has 300 milliliters (mL) of solution that is 18% acid. How many milliliters of acid are in the solution?

Answer by learning b(2)

a/5000=6.5/100=100a=6.5 times 5000
=32500/100 A =$325